write y in terms of x, if y is proportional to the 5th power of x, and y=744 when x=2
Answer by stanbon(75887)

write y in terms of x,
if y is proportional to the 5th power of x
y = kx^5
--------------
and y=744 when x=2
744 = k2^5
744 = k*32
k = 23.25
----------------
EQUATION:
y = 23.25x^5
=====================
Cheers,
Stan H.